OpenStudy (anonymous):

A 3.5 oz. serving of pudding is to be served to each of the 185 people at a dinner. How many cans must be ordered if each can contains 2lb 8 oz of pudding

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (anonymous):

So first you need to find how much pudding is needed to serve all 185 people. 3.5 * 185

OpenStudy (anonymous):

703

OpenStudy (anonymous):

@katlin95 u still there

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Not quite, 3.5 * 185 = 647.5

OpenStudy (anonymous):

lol! how did i get 702? lol. I made a mistake on calculator

OpenStudy (anonymous):

3.5⋅185÷2.8=231.25

OpenStudy (anonymous):

I think you typed 3.8 instead of 3.5

OpenStudy (anonymous):

katlin date me

OpenStudy (anonymous):

ur cute

OpenStudy (anonymous):

yea hah, so whats the next step? @sylvester13 im married, and have a child. thanks tho

OpenStudy (anonymous):

yo ur stupid bro lol ;)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

btw my picture is about 7 years old..

OpenStudy (anonymous):

oh that makes for sence

OpenStudy (anonymous):

@sylvester13 boy u look sick lol ;)

OpenStudy (anonymous):

3.5⋅185÷2.8=231.25 thats the best i can do for u ;) lol

OpenStudy (anonymous):

3.5⋅185÷2.8=231.25 thats the best i can do for u ;) lol dos that work @katlin95

OpenStudy (anonymous):

So now we need to find how many cans of pudding is needed to feed all 185 people. We know we need 647.5 oz of pudding and there are 2lb 8oz of pudding in each can. Can you find out how many total oz of pudding is in each can?

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

yes u multiply

OpenStudy (anonymous):

lol sorry dont listen to me im off track im rong

OpenStudy (anonymous):

we need to find how many total oz is in each can. (hint: 16oz = 1lb)

OpenStudy (anonymous):

32 oz.

OpenStudy (anonymous):

Almost :) 2 * 16 + 8 = ?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

Yes :) So now we know we need 647.5 oz of pudding and there are 40oz of pudding in each can. So how many cans are needed to get 647.5oz of pudding? Do not round your answer.

OpenStudy (anonymous):

17

OpenStudy (anonymous):

i got it now:) i checked the answer book and 17 is right

OpenStudy (anonymous):

You are just 0.8125 off :) 647.5 / 40 = 16.1875 So if you are rounding your answer it would be 16.

OpenStudy (anonymous):

Oh, ok :) I thought it was one off :\

OpenStudy (anonymous):

:) thanks!

OpenStudy (anonymous):

no prob :)